If alpha is the angle of limiting friction, then the minimum force required to pull a block of mass m kept on the horizontal surface is?


Ans mgsin(theta)

Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Nehal

Let a force F is applied on the block at an angel x from the horizontal

normal force will be N =mg - Fsinx

 so Fcosx =μN

     Fcosx =μ(mg - Fsinx)

   F =μmg/(cosx +μsinx)

for minim F   , (cosx +μsinx)  will be maximum so  (cosx +μsinx)=√(1+μ2)

So  F min = μmg/√(1+μ2)

here μ = tanΘ

so F min = tanΘmg/√(1+tan2Θ)

             =mgsinΘ

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin